OpenStudy (anonymous):

Can some one help me -(5a+6)=2(3a+8)

jimthompson5910 (jim_thompson5910):

-(5a+6)=2(3a+8) -5a-6=6a+16 -5a=6a+16+6 -5a-6a=16+6 -11a=16+6 -11a=22 a=22/(-11) a=-2